OpenStudy (anonymous):

find the complex zeros of the polynomial function. f(x)= 2x^4-11x^3+7x^2+149x-75 use complex zeros to write f in factored form

OpenStudy (anonymous):

using rational root theorem find some of ratioanl roots...then factorize the polynomial

OpenStudy (anonymous):

real roots are x=-3,1/2 complex roots are x=4+3i and x=4-3i